OpenStudy (anonymous):

can someone check my answers? simplify the following expression 21(x-4)+2 A.21x=2 my answer B. 107x C. 21x-82 D. 27x solve the following expression 5x-102=13 A. 23 my answer B. -23 C. -17.8 D. 17.8 solve the following equation 2x-31=3(x+2) A. 37 B.-37 my answer C. 5 D. 37/5 find the range of the function f(x)=4x=1 for the domain (-1,0,1,2,3) A. -5,-3,0,7,11 my answer find the slope of the graph 9x-3y=15 A: -3 B.3 my answer C. -1/3 D.1/3 if i got it wrong walk me through it and we will figure it out together:) @Nnesha @jim_thompson5910

OpenStudy (anonymous):

i can give the rest of the options for the range function if i got that wrong

OpenStudy (hockeychick23):

the first one is incorrect, you have to distribute the 21 so do 21*-4 and 21*x then last add 2

OpenStudy (anonymous):

#2 is C

Nnesha (nnesha):

for first one you have to distribute parenthesis by 21 2nd) right 3rd) right 4th one i guess there is a typo 4x plus or negative 1 ??

Nnesha (nnesha):

last one is right :)

OpenStudy (anonymous):

sorr its -1

OpenStudy (hockeychick23):

no the 2nd one you got right, solve the following expression 5x-102=13 add 102 to both sides. 5x-102+102=13+102 5x= 115 Divide both sides by 5 5x/5=115/5 x=23

Nnesha (nnesha):

okay  domain is all x values so replace x by ( (-1,0,1,2,3) ) to find range so when you replace x by 0 whats your answer suppose to be ??

OpenStudy (anonymous):

ok is number 1 21x-82? thats what i ended up with next

Nnesha (nnesha):

yes right :)

OpenStudy (anonymous):

-4?

Nnesha (nnesha):

4 times 0 = ?? and then subtract 1

OpenStudy (anonymous):

-3

Nnesha (nnesha):

4 times 0 = ?? its not equal to 4 when you multiply something by 0 you should get 0

OpenStudy (anonymous):

RIGHT SORRY i must of been thinking of 4 x 1

Nnesha (nnesha):

yep its okay :)

OpenStudy (anonymous):

of so we have 0 now

Nnesha (nnesha):

4 times 0 = 0 but you have to subtract one right 0 - 1 = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

-1

Nnesha (nnesha):

yep u got it :) -1

Nnesha (nnesha):

now replace x by 1 to find range

OpenStudy (anonymous):

by the 1 or -1?

Nnesha (nnesha):

 you already replaced x by -1 now by just one they gave you alll x values you have to plug that into x one by one to get range